What is specific gas constant? Is it same for all Gases?
 

Grade:12

6 Answers

venky
8 years ago
The Individual Gas Constant - R
Gas Individual Gas Constant - R  
Gas Imperial Units (ft lb/slugoR) SI Units (J/kg K)
Methane - natural gas, CH4 3099 518.3
Nitrogen, N2 1775 296.8
Oxygen, O2 1554 259.8
HARINATH
8 years ago
i) For 1 gm of a gas
PV
=r.
T
For m grams of the gas
PV
= mr ‘r’ is the specific gas
T
constant. ii) The specific gas constant is different for different gases.
 
HARINATH
8 years ago
i) For 1 gm of a gas
PV
=r.
T
For m grams of the gas
PV
= mr ‘r’ is the specific gas
T
constant. ii) The specific gas constant is different for different gases.
 
sathish
8 years ago
) For 1 gm of a gas PV =r. T For m grams of the gas PV = mr ‘r’ is the specific gas T constant. ii) The specific gas constant is different for different gases.
Surendra
8 years ago
The maximum height of mountain on earth depends upon shear modulus of rock. At thebase of the mountain, the stress due to all the rock on the top should be less than thecritical shear stress at which the rock begins to flow. Suppose the height of the mountainis h and the density of its rock is ρ . Then force per unit area (due to the weight of themountain) at the base = hρ gThe material at the experience this force per unit area in the vertical direction, but sidesof the mountain area free. Hence there is a tangential shear of the order of hρ g . Theelastic limit for a typical rock is about 3 ×108 Nm −2 and its density is 3 ×103 kgm −3Hence hmax ρ g = 3 ×108Or hmax3 × 1083 × 108=== 10, 000m = 10kmρg3 × 103 × 9.8This is more than the height of the Mount Everest.
ARATI SREEKANTH
8 years ago
 
i) For 1 gm of a gas
PV
=r.
T
For m grams of the gas
PV
= mr ‘r’ is the specific gas
T
constant. ii) The specific gas constant is different for different gases.